OpenStudy (anonymous):

I am trying to find the points on the line y=x+2 that is closet to (2, 1). Do not tell me the answer, but give me a head start on how I should begin

OpenStudy (anonymous):

READY SET GO!

OpenStudy (phi):

you probably want the line perpendicular to the given line that goes through pony (2,1)

OpenStudy (ddcamp):

The closest point will be on a line perpendicular to the original line that passes through the given point: 

OpenStudy (anonymous):

so I find the line perpendicular to the equation above

OpenStudy (ddcamp):

There's lots of lines that are perpendicular to that line:  You find the perpendicular slope, then use the point-slope equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Sorry if I am asking a lot of question, so I find the perpendicular slope, then the point-slope equation to give me a new equation?

OpenStudy (ddcamp):

Yes.

OpenStudy (ddcamp):

The point will be the intersection of the two equations.

OpenStudy (anonymous):

I did it and it does not make any sense. I found the perpendicular slope which is -1 and I found the new line equation which is -x+3. how does this give me the points. I plugged 2 as my x value and it gave me 1 as my y value.

OpenStudy (ddcamp):

y=-x+3 intersects y=x+2: -x+3 = x+2 -1 = -2x x = 1/2 y = -(1/2) + 3 y = 2+1/2

OpenStudy (anonymous):

so the point is (.5, 1.5). thanks a lot. what does it mean to approximate percent change

OpenStudy (ddcamp):

Sorry, not a clue what approximate percent change means.

OpenStudy (anonymous):

ok thanks, u've done more than enough.

OpenStudy (phi):

what is the original question ?

OpenStudy (anonymous):

approximate change in area of equilateral triangle if side length changes from 5 to5.1. what is the approximate percent change. I know how to do the first part. but not the second one

OpenStudy (phi):

write down the area of the triangle as a function of x (side length)= f(x) take the derivative of f(x) use the approximation formula \[ \Delta A = f' \cdot \Delta s \] percent change will be \[ \frac{\Delta A}{A} \] changed to a percent

OpenStudy (phi):

what did you get for \( \Delta A\) ?

OpenStudy (phi):

to be more clear, use \[ \Delta A = f'(5) \Delta x \] where x is the side

OpenStudy (anonymous):

got it. thanks

OpenStudy (anonymous):

I need help with this problem. An athelete who is 5'6" walks away from a lamp stand that is 22 foot tall at the rate of 3 feet per seconds. at what rate is the top of her shadow moving when she is 20 feet from the lampstand

OpenStudy (phi):

Have you seen http://tutorial.math.lamar.edu/Classes/CalcI/RelatedRates.aspx ? He gives an example of most types of related rates problems. See example 6